Can you please help me set up this proportional equation?
x+9/7 = x-12/4

Answer by stanbon(75887)

(x+9)/7 = (x-12)/4
Cross-multiply to get:
4(x+9) = 7(x-12)
4x+36 = 7x-84
3x=120
x=40
Cheers,
Stan H.